# Discovering Knowledge-Critical Subnetworks in Pretrained Language Models

### v1

@@ 4.2 Mask Learning
Suppression Loss   To fulfill the suppression criterion, the remaining model, denoted as $f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$, should be less confident in the expression of knowledge in $K_{T}$. We propose to minimize the KL divergence between the remaining model’s predicted distribution over possible tail entities of a knowledge triplet and a uniform reference distribution $p_{u}$ over the tokens in the model’s vocabulary. Thus, for $x\in K_{T}$:
$\mathcal{L}_{\text{suppress}}=D_{\text{KL}}(p_{u}\;\|\;f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}}))$ (3)
Maintenance Losses   As there are multiple ways a model could learn to suppress the expression $K_{T}$, mainly (1) suppressing all knowledge that is in the same format or (2) suppressing all language expressions completely, we define two regularization objectives. To encourage the rest of the model to keep its original performance on the control knowledge $K_{C}$ and a standard language modeling dataset $D_{LM}$, we calculate the KL divergence of $f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$ with the pretrained model’s distribution $f(x,{\bm{\theta}})$ as the reference. Therefore, for $x\in K_{C}$ and $x\in D_{LM}$ :
@@
Sparsity Regularization   To encourage our subnetwork to be sparse for maintenance reasons (i.e., reducing side effects to pretrained model behavior when removed) and so that they do not contain non-critical parameters for modeling TargetKG (e.g., redundant language modeling parameters), we minimize the average subnetwork density (i.e., sigmoid of the masking parameters ${\bm{l}}_{i}$ from Eq. 1):
@@
Final Loss   Our final loss is a mixture of these losses with weights $\lambda_{i}$ (listed in Appendix B):
$\mathcal{L}_{\text{final}}=\lambda_{1}\mathcal{L}_{\text{suppress}}+\lambda_{2}\mathcal{L}_{\text{maintain-KG}}+\lambda_{3}\mathcal{L}_{\text{maintain-LM}}+\lambda_{4}\mathcal{L}_{\text{sparsity}}$ (6)
@@ 5 Experimental Setup
Models & Training   To test whether our method can scale to various model sizes, we discover knowledge subnetwork masks for GPT2-small (117M parameters, 12 layers), GPT2-medium (345M parameters, 24 layers), and GPT2-large (774M parameters, 36 layers; Radford et al., 2019). During mask learning, we do not mask the embedding, language modeling head, layer-normalization, and bias parameters. [^1] Prior work has not observed an advantage to masking these components for general tasks (Zhao et al., 2020). We also only learn masks for the top 50% of the transformer layers. [^2] Multiple layer-wise model analyses have shown that the first few layers of transformer language models encode low-level linguistic tasks and features that may be a prerequisite for knowledge modeling (Tenney et al., 2019; Liu et al., 2019). We also perform a masked layer choice study that confirms this intuition (Appendix C). For more information on implementation and checkpoint selection, please refer to Appendix B.
@@
Datasets  To create TargetKG and ControlKGs, we sample hypernym triplets from WordNet (Miller, 1995), as well as triplets from the LAMA subset of ConceptNet (Speer et al., 2017; Petroni et al., 2019). For simplicity, we only consider triplets with single-token tail entities. To gather small connected TargetKG graphs, we randomly select an initial node and sample knowledge triplets by walking a depth of three up (parent direction) and down (child direction) in the respective KG. We sample 7 TargetKGs for WordNet using this method, and 3 for ConceptNet (statistics shown in Table 6 of the Appendix). To create ControlKG, we prioritize not leaking TargetKG counterfactuals and having a shared ControlKG across different TargetKGs, and so remove from the complete KG any triplet that shares the same entities as the union of the TargetKGs shown in Table 6. For all KG verbalizations, to remove and maintain knowledge that the model is already confident about, we pick the best scoring verbalization for each triplet among several prompt styles. Statistics on TargetKG and ControlKG datasets can be seen in Table 6. For the ControlLM dataset, we use WikiText-2 (Merity et al., 2017). We refer to ControlKG and ControlLM together as maintenance datasets. The ControlKG and ControlLM results are on the held-out validation set. Please refer to Appendix A and B for more data processing details and examples.
Success Metrics   Considering perplexity (PPL) as a proxy for a model’s confidence in the expression of knowledge, we can reformulate the knowledge-critical subnetwork goals as:
  1. 1.
    Suppression: $\text{PPL}(f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}}))\ll\text{PPL}(f(x,{\bm{\theta}}))$, for $x\in K_{T}$
  2. 2.
    Maintenance-KG: $\text{PPL}(f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}}))\approx\text{PPL}(f(x,{\bm{\theta}}))$, for $x\in K_{C}$
  3. 3.
    Maintenance-LM: $\text{PPL}(f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}}))\approx\text{PPL}(f(x,{\bm{\theta}}))$, for $x\in D_{LM}$
  4. 4.
    Sparsity: $0<\sum_{i=1}^{|{\bm{\theta}}|}{\bm{m}}_{i}\ll|{\bm{\theta}}|$
@@
To measure these conditions, we calculate the perplexity difference between the remaining and original models. We refer to the perplexity difference as $\Delta$ PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$). For the suppression and maintenance-KG criteria, we calculate $\Delta$ PPL using the loss on the masked tail entity for examples in the TargetKG and ControlKG datasets. For a knowledge-critical subnetwork, we expect $\Delta$ PPL to be high for TargetKG and low for ControlKG. For the maintenance-LM criterion, we calculate $\Delta$ PPL as the average perplexity on all tokens in a sequence, which should be low if the knowledge-critical subnetwork mask does not affect the model’s general language modeling ability. For the sparsity criterion, we calculate the percentage of mask parameters that are 0 after the straight-through threshold in Equation 2. The denominator is the number of maskable parameters. Ideally, the sparsity should be as large as possible (e.g., near 99%).
@@
Baseline   As a control baseline, we create randomly masked models at the same sparsity level as the knowledge-critical subnetwork. If the discovered subnetwork is critical for expressing TargetKG, then removing a random subnetwork at the same sparsity should have significantly lower corruption for expressing TargetKG (i.e., lower $\Delta$ PPL) than removing the critical subnetwork. Similarly, if the critical subnetwork successfully preserves the maintenance criteria, a random subnetwork should be more likely to remove useful weights for expressing ControlKG and ControlLM, which should lead to a higher $\Delta$ PPL on maintenance datasets. For more information on the baseline implementation, please refer to Appendix B.
@@ 6 Experimental Results
6.1 Knowledge-Critical Subnetwork Discovery

### v2

@@ 4.2 Mask Learning
Suppression Loss
To fulfill the suppression criterion, the remaining model, $f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$, should be less confident in the expression of knowledge in $K_{T}$. We propose to minimize the KL divergence between the remaining model’s predicted distribution over possible tail entities of a knowledge triplet and a uniform reference distribution $\mathcal{U_{V}}$ over the tokens in the model’s vocabulary. For $x\in K_{T}$:
$\mathcal{L}_{\text{suppress}}=D_{\text{KL}}(\mathcal{U_{V}}\;\|\;f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}}))$ (3)
Maintenance Losses
As there are multiple ways a model could learn to suppress the expression $K_{T}$, namely (1) suppressing all knowledge that is in the same format or (2) suppressing all language expressions, we define two regularization objectives. To encourage the rest of the model to keep its original performance on the control knowledge $K_{C}$ and a standard language modeling dataset $D_{LM}$, we calculate the KL divergence of $f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$ with the pretrained model’s distribution $f(x,{\bm{\theta}})$ as a reference. Thus, for any $x\in K_{C}$ or $x\in D_{LM}$:
@@
Sparsity Regularization
To promote the subnetwork containing only parameters critical for modeling TargetKG, we encourage sparsity by minimizing the average subnetwork density (i.e., sigmoid of the masking parameters ${\bm{l}}_{i}$ from Eq. 1):
@@
Final Loss Our final loss is a mixture of these losses with weights $\lambda_{i}$:
$\begin{split}\mathcal{L}_{\text{final}}=\lambda_{1}\mathcal{L}_{\text{suppress}}+\lambda_{2}\mathcal{L}_{\text{maintain-KG}}\\ +\lambda_{3}\mathcal{L}_{\text{maintain-LM}}+\lambda_{4}\mathcal{L}_{\text{sparsity}}\end{split}$ (6)
@@ 5 Experimental Setup
To test whether our method can scale to various model sizes, we discover knowledge subnetwork masks for GPT2-small, (117M parameters, 12 layers), GPT2-medium, (345M parameters, 24 layers), GPT2-large, (774M parameters, 36 layers), and GPT2-XL. (1.5B parameters, 42 layers; Radford et al., 2019). During mask learning, we do not mask the embedding, language modeling head, layer-normalization, and bias parameters, [^2] Prior work has not observed an advantage to masking these components for general tasks Zhao et al. (2020). and only learn masks for the top 50% of transformer layers. [^3] Multiple layer-wise analyses have shown that the first layers of transformer LMs encode low-level linguistic features that may be a prerequisite for knowledge modeling Tenney et al. (2019); Liu et al. (2019a). We also perform a masked layer choice study that confirms this intuition (Appendix C). Further implementation details on masking, hyperparameter, and checkpoint selection are in Appendix B.
Datasets
@@
To create TargetKG and ControlKGs, we sample hypernym triplets from WordNet Miller (1995), as well as triplets from the LAMA subset of ConceptNet Speer et al. (2017); Petroni et al. (2019). For simplicity, we only use triplets with single-token tail entities. We sample 7 TargetKGs for WordNet, and 3 for ConceptNet (statistics shown in Table 1) by randomly selecting an initial node and sampling knowledge triplets by performing 3-hop random walks in both the parent and child direction of the KG. To create ControlKG, we prioritize not leaking TargetKG counterfactuals and having a shared ControlKG across different TargetKGs, and remove from the complete KG any triplet that shares the same entities as the union of the TargetKGs shown in Table 1. For all triplets, to suppress and maintain knowledge that the model is already confident about, we select the verbalization for each triplet with the lowest perplexity on the tail token. For the ControlLM dataset, we use WikiText-2 Merity et al. (2017). We refer to ControlKG and ControlLM together as maintenance datasets. All maintenance results are on a held-out validation set. Further information on data preprocessing is in Appendices A and B.
Knowledge Mask Sparsity TargetKG ControlKG ControlLM TargetKG ControlKG
Graph Method ($\uparrow$) $\Delta$PPL ($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$) $\Delta$Rank ($\uparrow$) $\Delta$Rank ($\downarrow$)
WordNet Weight Masking 98.6 590.9 (162.4) -0.2 (73.2) 0.5 (37.7) 320.9 (45.4) 1.4 (9.7)
Neuron Masking 95.3 715.9 22.2 4.1 288.7 7.1
Random Weights 98.6 24.3 14.6 2.2 12.0 2.7
Random Neurons 95.3 23.8 8.3 8.3 17.0 4.5
ConceptNet Weight Masking 99.1 636.4 (429.0) 2.8 (172.3) 0.2 (37.7) 636.1 (290.6) 1.6 (47.5)
Neuron Masking 94.9 22422.0 71.9 5.4 8720.2 29.4
Random Weights 99.1 21.0 14.6 1.5 11.4 5.5
Random Neurons 94.9 110.7 70.4 11.2 35.9 28.5
Table 2: Subnetwork discovery for GPT2-small, averaged over three seeds and seven KGs for WordNet, and three KGs for ConceptNet. $\Delta$PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$) and similarly for $\Delta$Rank results. The values in parenthesis are the average metric (PPL or Rank) for the pretrained model (i.e., the base from which the $\Delta$ is computed). The arrows ($\uparrow$,$\downarrow$) show the desired direction for the metric. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair.
Metrics
@@
We follow prior work Hase et al. (2023a) that considers perplexity (PPL) as a proxy for a model’s confidence in the expression of knowledge, and calculate the perplexity difference between the remaining and original models, $\Delta$PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$). We also report $\Delta$Rank, the tail token rank difference between the remaining and original models. For the suppression and maintenance-KG criteria, we calculate $\Delta$PPL using the loss on the masked tail entity for triplets in the TargetKG and ControlKG datasets. For a knowledge-critical subnetwork, we expect $\Delta$PPL and $\Delta$Rank values to be high for TargetKG and low for ControlKG. For the maintenance-LM criterion, we calculate $\Delta$PPL as the average perplexity on all tokens in a sequence, which should be low if removing the critical subnetwork does not affect the model’s general language modeling ability. [^4] We do not report $\Delta$Rank for maintenance-LM as the average rank of all tokens in an open-ended sentence is not as informative as the single tail token rank. For the sparsity criterion, we calculate the percentage of parameters that were not pruned. The denominator is the number of masked parameters, meaning the total size of dense layers in the upper half of the model. Ideally, the sparsity should be as high as possible to keep the majority of parameters (i.e., near 99%).
Baseline
@@
We use weight and neuron masking to localize knowledge-critical subnetworks. As a control baseline, we create randomly masked models at the same sparsity level as the knowledge-critical subnetwork. If the discovered subnetwork is critical for expressing TargetKG, then removing a random subnetwork at the same weight or neuron sparsity should yield lower corruption for expressing TargetKG (i.e., lower $\Delta$PPL) than removing the critical subnetwork. Similarly, if the critical subnetwork successfully preserves the maintenance criteria, a random subnetwork should be more likely to prune useful weights for expressing ControlKG and ControlLM, which should lead to a higher $\Delta$PPL on maintenance datasets. Further information on the implementation of the random masking baseline is in Appendix B.
@@ 6 Experimental Results
We first evaluate the degree to which discovered subnetworks are knowledge-critical.
Ablation Sparsity TargetKG ControlKG ControlLM
($\uparrow$) $\Delta$PPL ($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
Table 3: Ablation study for the multi-objective loss on GPT2-small using weight masking, with [min, max] boundaries, averaged across three KGs and two seeds.
Weight-masked Subnetworks
In Table 2, we observe that across seven different knowledge graphs (TargetKGs) and three random seeds, the subnetworks found with weight masking consistently achieve a notably high sparsity ($>$ 98%). [^5] Table 13 provides individual KG results for the averaged weight masking results in Table 2. For the suppression criterion, we notice a high $\Delta$PPL on TargetKG for both approaches, meaning that the perplexity of the remaining model on TargetKG is significantly higher than the pretrained model’s perplexity. In contrast, removing a random subnetwork at the same sparsity yields a smaller perplexity increase, meaning the discovered subnetworks are significantly more critical for expressing TargetKG. At the same time, we find little change in perplexity on the maintenance datasets for relational knowledge (ControlKG) and language modeling (ControlLM), demonstrated by the negligible $\Delta$PPL on both datasets and the small $\Delta$Rank value on ControlKG.  [^6] Note that the lower average PPL of ControlKG compared to TargetKG is due to ControlKG being larger, which minimizes the impact of outliers and reduces average perplexity. We note that a negative $\Delta$PPL here may result from the remaining model slightly overfitting to the ControlKG distribution, although it is never too significant.
We observe similar results for knowledge-critical subnetworks for larger models. For three TargetKGs: communication, representation, and location, we observe an average increase in TargetKG perplexity of 256 for GPT2-medium, 5780 for GPT2-large, 536 for GPT2-XL, and a negligible maintenance $\Delta$PPL (Table 16).
Neuron-masked Subnetworks
On the other hand, neuron masking does not reliably fulfill the conditions of discovering knowledge-critical subnetworks. While removing neuron-masked subnetworks yields greater suppression of TargetKG than weight masking, it also significantly impacts ControlKG $\Delta$PPL and $\Delta$Rank (more than randomly removing neurons at the same sparsity), indicating that other behaviors of the model are not robustly maintained. They also tend to be less sparse, frequently keeping $\sim$5of the parameters of the original model. [^7] Appendix Table 14 provides individual KG results for the averaged neuron masking results in Table 2. We hypothesize that this observation is potentially related to neuron superposition (Elhage et al., 2022), where the neurons that represent TargetKG cannot be fully disentangled from representations that encode general relational knowledge. While weights may also be polysemantic, they are more fine-grained, potentially encoding knowledge in a more separable manner.
Ablation Study
As our method relies on a joint objective combining multiple loss functions, we perform an ablation study of the loss terms presented in §4.2 for weight masking and remove each objective (i.e., No Suppression, No Maintenance-KG, No Maintenance-LM) to validate whether these losses accomplish their goals. [^8] We do not ablate the sparsity term. Without it, the subnetwork search stagnates at the initial sparsity. In Table 3, we observe that the suppression loss is necessary to increase TargetKG perplexity (and suppress the knowledge). Without it, the model only optimizes for retaining ControlKG, and generalizes this improvement to TargetKG as well (as indicated by the negative $\Delta$PPL). We also find that removing the maintenance losses significantly affects ControlKG and ControlLM perplexity differences. Without these controls, our method learns to suppress the knowledge from the model by suppressing general abilities. The suppression objective, a minimization of the KL divergence between the output distribution and a uniform distribution, affects the prediction of tail entities for all relational knowledge rather than affecting only TargetKG. We present additional ablations related to varying the training objectives in Appendices B (varying $\lambda_{i}$ in Eq. 6) and F (adding additional loss terms).
Paraphrase Generalization
To assess whether our subnetworks generalize to other verbalizations of TargetKG and ControlKG, we evaluate the pruned models on 20 other distinct relation paraphrases that are not used during training. Specifically, we vary the tokens representing the relation and the format of the head and tail entities while still ensuring grammatical correctness. [^9] For further details and examples on the creation of these verbalizations, please refer to Appendix A. For weight masking, our conclusions do not change when using other prompt styles, as seen in Table 4. Interestingly, the $\Delta$PPL for ControlKG paraphrases is sometimes lower than for the format used for training, likely because the starting perplexity is higher on other templates, [^10] Recall that for every triplet, we use the verbalization with the lowest original model perplexity for training. The average perplexity of GPT2-small on the worse paraphrases is 3231 for TargetKG and 2368 for ControlKG. and the maintenance of ControlKG generalizes to a greater degree on these suboptimal templates. The neuron masking approach generalizes well to TargetKG paraphrases, but poorly for ControlKG templates, reinforcing our previous observations.
Mask Knowledge TargetKG ControlKG
Method Graph $\Delta$PPL ($\uparrow$) $\Delta$PPL ($\downarrow$)
Weight communication 492.4 3.0
location 684.9 -33.0
representation 916.4 -6.9
Neuron communication 866.8 -325.0
location 941.6 284.5
representation 1928.9 654.5
Table 4: Paraphrase results on GPT2-small.
6.1 Subnetwork Analysis
Subnetwork Structure
To better understand how knowledge-critical subnetworks interact with the rest of the model, we explore their structure in the parameter space of the original model. For three WordNet TargetKGs and three random seeds, we find that GPT2-small subnetworks are relatively denser in the first and final masked transformer blocks. For weight masking, more density is observed in the attention sublayers (Figure 3). Interestingly, much of the density of the subnetworks in the attention sublayers is tied to individual attention heads (Figure 5), supporting prior conclusions that particular attention heads encode semantic relationships (Clark et al., 2019; Geva et al., 2023).
However, despite being dense around similar portions of the model across different TargetKGs and random seeds, the subnetworks are quite distinct. When we calculate the Jaccard similarity (i.e., IoU) of the individual parameters across subnetworks for different random seeds for the same TargetKG, the result is quite low on average for weight-masked subnetworks (3-4%) — though higher for the final attention output sublayer (10-12%) — indicating the knowledge-critical subnetworks are quite disjoint, even when discovered by suppressing the same information (Figure  10).
Neuron masking led to a much higher density in the second feedforward layers of the transformer blocks and attention layers (Figure 4). We find that the IoU of neuron-masked subnetworks are also 10$\times$ higher (34-44%; Figure 11), partially due to their reduced sparsity, but also perhaps indicating that neuron masking yields more unique subnetworks across seeds, though they are also less reliably knowledge-critical.
Mask Composition Sparsity TargetKG ControlKG ControlLM
Pattern ($\uparrow$) $\Delta$PPL ($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
Weight Individual 98.8 379.1 0.7 0.4
Union 96.9 1984.4 44.9 4.7
Floral 99.5 4.9 4.5 1.1
Intersection 99.9 7.9 3.7 2.1
Neuron Individual 95.2 396.0 22.3 4.1
Union 92.8 2743.0 86.9 6.6
Floral 95.5 290.5 16.1 3.7
Intersection 97.3 63.9 3.5 2.4
Table 5: Composing subnetworks with GPT2-small. Individual stands for the individual subnetwork removal average across the same three seeds and KGs.
Subnetwork Composition
However, even though knowledge-critical subnetworks across random seeds may be disentangled, composing them (and removing them jointly) amplifies the suppression effect. As shown in Table 5, when we compose subnetworks for GPT2-small as a union of three random seed masks for the same TargetKG, the suppression effect increases significantly, by a factor of 6$\times$ (far more than removing additional random parameters from the remaining model; Figure 2). While this suppression is accompanied by a degradation in the maintenance criteria ($\sim$30-40 $\Delta$PPL on ControlKG instead of near 0), the absolute difference is far smaller. Composing neuron-masked subnetworks yields similar trends, though we observe two interesting patterns. First, the intersection of these subnetworks produces a subnetwork that satisfies the maintenance criteria to be knowledge-critical, though at the cost of reducing suppression. Second, neuron-masked compositions yield monotonic changes in suppression and maintenance scores as sparser composition methods are used. Further analyses on seed-based and knowledge-based variance across discovered subnetworks are in Appendix I and J, respectively.
Subnetwork Sensitivity
Finally, we investigate whether discovered subnetworks are structurally sensitive. Specifically, we perform a sensitivity analysis of the recorded metrics as we iteratively expand or contract the subnetwork (by adding or removing parameters). As we add parameters to the subnetwork (i.e., remove parameters from the remaining model), we measure the change in TargetKG $\Delta$PPL. In this case, a sudden drop in $\Delta$PPL would indicate that the discovered subnetwork is spurious. In Appendix Figure 2, we observe that expanding the discovered subnetwork in small amounts does not significantly recover the model’s ability to express TargetKG, providing further evidence that the subnetworks are not arbitrarily discovered, but rather have meaningful knowledge-expressing structure within the larger model. We provide more experimental details in Appendix G.
6.2 Downstream Task Transfer
If a subnetwork is truly knowledge-critical, its removal should harm a pretrained language model’s ability to transfer to a downstream task requiring the knowledge encoded by the subnetwork. To test this hypothesis, we finetune a model on the CommonsenseQA benchmark Talmor et al. (2019) after removing a relevant knowledge-critical subnetwork. We use the in-house splits from Lin et al. (2019), with a development set of 1241 and an initial test set of 1221 questions. In the test set, we induce [^11] We describe this process in Appendix E. the ConceptNet relation linked to each question and extract the relevant triplets from ConceptNet, creating a TargetKG from all ConceptNet triplets associated to the test set, which yields a filtered set of 363 questions for which we can reliably extract relevant ConceptNet triplets. We use these relevant triplets as TargetKG and the remaining distinct triplets in the LAMA subset of ConceptNet as ControlKG to learn a knowledge-critical subnetwork using either weight and neuron masking for GPT2-small. Then, we apply different finetuning methods to the remaining model after removing the critical subnetwork, using the same training set. We compare finetuning the remaining masked model (Weight Mask, Neuron Mask in Table 6) to the performance of finetuning the full pretrained model (Full), as well as a randomly masked model at the same sparsity as the masked-weight subnetwork (Random Mask). We report results across three random seeds in Table 6.
For all finetuning methods, we find that the remaining model with weight masking has similar accuracy to the pretrained model on the development split and a close accuracy for the overall test set. However, we observe a consistent significant performance drop on the filtered subset after finetuning (average drop of 7.3%; head tuning barely better than selecting a random answer on a 5-choice MCQA task), indicating that the model struggles to transfer knowledge associated with TargetKG during fine-tuning. Interestingly, in less parameter-efficient finetuning methods, this drop does not persist when the neuron-masked subnetwork is removed, suggesting that knowledge is either still transferred or recovered over the course of finetuning (Lo et al., 2024). In addition, for both head tuning and LoRA Hu et al. (2022) with weight masking, we find that if we randomly split the filtered TargetKG, one half’s knowledge-critical mask does not affect the accuracy of the other half as significantly as its own (see Appendix E for details), indicating the performance drop is indeed specific to the pruned knowledge.
Tuning Method Subnetwork Dev Test Filtered
Head Tuning Full 38.63 38.33 37.19
Random Mask -0.47 -1.61 -3.21
Neuron Mask -3.74 -4.22 -6.61
Weight Mask -1.69 -6.80 -14.42
LoRA Full 50.04 48.64 48.67
Random Mask -0.74 -2.33 -1.75
Neuron Mask -0.30 +1.89 +1.28
Weight Mask -1.83 -2.74 -3.95
Full Finetuning Full 44.61 42.33 42.79
Random Mask +0.30 -0.24 +2.39
Neuron Mask +0.38 +2.17 +1.47
Weight Mask -1.50 -5.14 -3.60
Table 6: Accuracy on CommonsenseQA, averaged over three seeds for GPT2-small.
7 Conclusion
In this paper, we conceptualize knowledge-critical subnetworks, sparse computational subgraphs within larger language models that are responsible for expressing specific knowledge relationships. We discover these subnetworks using a multi-objective differentiable masking approach that jointly optimizes a criterion designed to suppress the expression of target knowledge when knowledge-critical subnetworks are removed from a language model, and maintenance criteria that ensure the language model retains its initial capacity to model other relational knowledge and general language. Our results show that when knowledge-critical subnetworks are removed, a model loses its ability to express the knowledge encoded in the subnetwork, and to transfer it when finetuned on downstream tasks requiring the knowledge.
Acknowledgements
We thank Mohammadreza Banaei, Syrielle Montariol, Debjit Paul, Khai Loong Aw, Badr AlKhamissi, Silin Gao, Yifan Hou, Beatriz Borges, Yu Fei, and Angelika Romanou for their helpful discussions and feedback on our manuscript. We also gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI.
Limitations
We discuss the limitations of our proposed method and conducted experiments on three axes: data, model, and hyperparameter. We emphasize that the data used for our experiments are limited to English only. As English is a high-resource language, additional challenges could arise when reproducing our method in a low-resource language (e.g., finding a rich lexical database like WordNet). We identify the lack of diverse pretrained language model architectures and language modeling objectives as the main model limitation. We have tested our method on the billion scale but did not expand our scope to larger models with different architectures (for example, in the 7B scale). We also limit the analysis to models trained with an autoregressive language modeling objective in contrast to text-to-text models such as T5 Raffel et al. (2020) or Masked-Language-Modeling models such as RoBERTa Liu et al. (2019b). Finally, the hyperparameter search detailed in the Appendix, while not exhaustive, provides sufficient evidence to support the validity of the selected range. To find more precise knowledge-critical subnetworks, future methods may need to take this hyperparameter search further.
Ethics Statement
In this study, we concentrate on relational knowledge, but the technique of identifying subnetworks could be used in mitigating bias within models. Likewise, this method of finding subnetworks may also inadvertently lead to the elimination of critical ethical or factual knowledge from a language model, resulting in a model that could generate offensive content and misinformation. For example, there exists a backdoor attack method against deep neural networks that builds on top of the identification and editing of subnetworks Qi et al. (2021). Therefore, caution should be exercised when applying the identification and removal of subnetworks to models used in essential applications.
Appendix A Dataset Creation and Processing
TargetKG
To gather small connected TargetKGs, we randomly select an initial node and sample knowledge triplets by walking a depth of three up (parent direction) and down (child direction) in the respective KG. Given a seed node such as representation in WordNet [^12] In WordNet, a word sense is represented by its lemma, syntactic category, and sense ID (e.g., in map.n.01, n for noun and 01 for sense ID). We omit this naming convention from the main paper tables for readability. or fruit in ConceptNet, we sample relations by performing a 3-hop random walk. For example, for the fruit KG shown in Table 7, we start from the seed concept fruit. In the first depth, we retrieve (fruit, ReceivesAction, eaten) and (wine, MadeOf, fruit). In the next depth, we retrieve (champagne, IsA, wine), and so forth for all possible relations. Note that we only sample relations with a single-token tail entity.
Finally, we verbalize TargetKG graph with the formats that give the lowest perplexity on the pretrained model. We try various relation-specific verbalization templates per knowledge triplet and pick the one that yields the lowest tail-token perplexity. For example, in the representation graph, while the model had lower perplexity with the template ‘‘{$h$} is a kind of {$t$}’’ for the triplet (representation.n.02, IsA, creation.n.02), it also had lower perplexity with the template ‘‘A {$h$} is a {$t$}’’ for the triplet (chart.n.02, IsA, map.n.01). Note that this can change for each model size, such as GPT2-small, medium, large and XL.
ControlKG
To create ControlKG, we prioritize not leaking TargetKG counterfactuals and having a shared ControlKG across different TargetKGs. Therefore, we remove from the complete KG (e.g., for ConceptNet TargetKGs, the complete LAMA subset of ConceptNet) any triplet that shares the same entities as the union of the TargetKGs shown in Table 1. For all KG verbalizations, to remove and maintain knowledge that the model is already confident about, we pick the best scoring verbalization for each triplet among several prompt styles and filter out those that yield an individual PPL higher than a threshold. For testing, we use held-out triplets.
ControlLM
We use WikiText-2 Merity et al. (2017) for the ControlLM dataset. We tokenize each entry and then concatenate all of them together. Finally, we group the tokens into chunks of 512. For validation and testing, we use held-out sets.
Appendix B Training and Evaluation Implementation
Init. Mask Sparsity TargetKG ControlKG ControlLM
Probability ($\uparrow$) $\Delta$PPL($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
0.25 99.5 [99.5, 99.6] 332.1 [119.5, 544.6] 3.7 [2.6, 4.7] 0.1 [0.1, 0.1]
0.45 99.5 [99.5, 99.5] 1287.8 [80.0, 2495.6] 0.9 [-1.1, 2.9] 0.2 [0.1, 0.2]
0.50 99.4 [99.4, 99.5] 939.1 [59.9, 1818.2] -0.2 [-0.2, -0.1] 0.2 [0.2, 0.2]
0.75 98.9 [98.8, 99.1] 13043.1 [4.6, 26081.5] -1.9 [-1.9, -1.8] 0.3 [0.3, 0.4]
Table 8: Hyperparameter study on inital mask probability with GPT2-small.
Masking $\bm{\lambda_{1}}$ $\bm{\lambda_{2}}$ $\bm{\lambda_{3}}$ $\bm{\lambda_{4}}$ Sparsity TargetKG ControlKG ControlLM Number of Valid
Method ($\uparrow$) $\Delta$PPL($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$) Checkpoints
Weight 1 1 3 1 96.9 [95.6, 98.2] 17.3 [-7.8, 42.4] 1.0 [-2.0, 4.1] 0.5 [0.2, 0.8] 2.5 [0, 5]
1 3 1 1 97.3 [96.4, 98.3] 80.9 [59.2, 102.6] -1.5 [-8.9, 5.8] 1.0 [0.5, 1.5] 53.5 [5, 102]
3 1 1 1 98.5 [97.6, 99.3] 12516.4 [682.2, 24350.6] 0.0 [-1.4, 1.5] 0.4 [0.2, 0.6] 145.0 [133, 157]
Neuron 1 1 3 1 96.7 [96.6, 96.9] 297.5 [104.0, 491.0] 3.9 [-1.7, 9.6] 2.9 [2.9, 3.0] 0.0 [0, 0]
1 3 1 1 95.3 [95.3, 95.4] 110.2 [53.3, 167.2] 11.8 [10.0, 13.7] 3.5 [3.4, 3.6] 0.0 [0, 0]
3 1 1 1 93.3 [93.1, 93.5] 255.1 [135.0, 375.1] 44.7 [38.8, 50.7] 5.2 [4.7, 5.7] 0.0 [0, 0]
Table 9: Hyperparameter study on $\lambda_{i}$ loss weights in Eq. 6 with GPT2-small.
Setup Model KG ControlLM
GPT-2 Small 250 10
Train GPT-2 Medium 96 4
GPT-2 Large 96 4
GPT-2 XL 128 4
GPT-2 Small 250 8
Eval GPT-2 Medium 250 8
GPT-2 Large 250 8
GPT-2 XL 250 8
Table 10: GPU batch size for each dataset and model.
Iteration TargetKG ControlKG ControlLM
$\Delta$PPL Floor $\Delta$PPL Ceiling $\Delta$PPL Ceiling
1 35.0 5.0 1.0
2 40.0 7.0 2.0
3 40.0 10.0 3.0
4 50.0 15.0 4.0
Table 11: Selection limit for each success criteria.
Masked Layer Percentage Sparsity TargetKG ControlKG ControlLM # of
Choice Masked ($\uparrow$) $\Delta$PPL($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$) checkpoints